# StreamMind: Unlocking Full Frame Rate Streaming Video Dialogue through Event-Gated Cognition

### v2

@@ 3 Experiment
| Method | Not use HT 100M | COIN Benchmark Top-1 Accuracy | ||||
| Step | Task | Next | Proc | Proc.+ | ||
| ClipBERT | ✓ | 30.8 | 65.4 | - | - | - |
| TimeSformer | 46.5 | 85.3 | 34.0 | 17.0 | 40.1 | |
| Paprika | 51.0 | 85.8 | 43.2 | - | - | |
| DistantSup | 54.1 | 90.0 | 39.4 | - | - | |
| VideoTF | 56.5 | 91.0 | 42.4 | 40.2 | 46.4 | |
| ProcedureVRL | 56.9 | 90.8 | 46.8 | - | - | |
| VideoTaskGraph | 57.2 | 90.5 | 40.2 | - | - | |
| VideoLLM-online-7B-v1 | ✓ | 59.8 | 92.1 | 48.1 | 47.9 | 52.9 |
| VideoLLM-online -8B-v1+ | ✓ | 63.1 | 92.7 | 49.1 | 49.8 | 54.1 |
| VideoLLM-MOD | ✓ | 63.4 | 92.8 | 49.7 | 49.8 | 53.3 |
| Ours | ✓ | 63.7 | 93.2 | 49.9 | 49.8 | 54.2 |
| Method | Not use Ego VLP | End-to -End? | Ego4D LTA ED@Z=20 | ||
| Verb | Noun | Action | |||
| CLIP | ✓ | ✓ | 0.739 | 0.769 | 0.941 |
| EgoT2 | ✓ | ✓ | 0.722 | 0.764 | 0.935 |
| I-CVAE | ✓ | ✓ | 0.753 | 0.749 | 0.931 |
| HierVL | ✓ | ✓ | 0.724 | 0.735 | 0.928 |
| VideoLLM | ✓ | 0.721 | 0.725 | 0.921 | |
| VideoLLM-online-7B-v1 | ✓ | ✓ | 0.697 | 0.698 | 0.897 |
| VideoLLM-online-8B-v1+ | ✓ | ✓ | 0.689 | 0.671 | 0.884 |
| VideoLLM-MOD | ✓ | ✓ | 0.689 | 0.676 | 0.884 |
| Ours | ✓ | ✓ | 0.689 | 0.655 | 0.881 |

### v3

@@ 3 Experiment
| Method |
COIN Benchmark Top-1 Accuracy | ||||||
| Step | Task | Next | Proc | Proc.+ | ||||
| ClipBERT | ✓ | 30.8 | 65.4 | - | - | - | ||
| TimeSformer | 46.5 | 85.3 | 34.0 | 17.0 | 40.1 | |||
| Paprika | 51.0 | 85.8 | 43.2 | - | - | |||
| DistantSup | 54.1 | 90.0 | 39.4 | - | - | |||
| VideoTF | 56.5 | 91.0 | 42.4 | 40.2 | 46.4 | |||
| ProcedureVRL | 56.9 | 90.8 | 46.8 | - | - | |||
| VideoTaskGraph | 57.2 | 90.5 | 40.2 | - | - | |||
| VideoLLM-online-7B-v1 | ✓ | 59.8 | 92.1 | 48.1 | 47.9 | 52.9 | ||
| VideoLLM-online -8B-v1+ | ✓ | 63.1 | 92.7 | 49.1 | 49.8 | 54.1 | ||
| VideoLLM-MOD | ✓ | 63.4 | 92.8 | 49.7 | 49.8 | 53.3 | ||
| Ours | ✓ | 63.7 | 93.2 | 49.9 | 49.8 | 54.2 | ||
| Method |
Ego4D LTA ED@Z=20 | ||||||
| Verb | Noun | Action | |||||||
| CLIP | ✓ | ✓ | 0.739 | 0.769 | 0.941 | ||||
| EgoT2 | ✓ | ✓ | 0.722 | 0.764 | 0.935 | ||||
| I-CVAE | ✓ | ✓ | 0.753 | 0.749 | 0.931 | ||||
| HierVL | ✓ | ✓ | 0.724 | 0.735 | 0.928 | ||||
| VideoLLM | ✓ | 0.721 | 0.725 | 0.921 | |||||
| VideoLLM-online-7B-v1 | ✓ | ✓ | 0.697 | 0.698 | 0.897 | ||||
| VideoLLM-online-8B-v1+ | ✓ | ✓ | 0.689 | 0.671 | 0.884 | ||||
| VideoLLM-MOD | ✓ | ✓ | 0.689 | 0.676 | 0.884 | ||||
| Ours | ✓ | ✓ | 0.689 | 0.655 | 0.881 | ||||